OpenStudy (anonymous):

Compute and simplify the difference quotient f(a+h)-f(a) --------- h g(x)=2x^2+14x -13

OpenStudy (anonymous):

this was already confusing enough without changing it to g(x)

OpenStudy (anonymous):

replace \(g\) by \(f\) it makes no difference, it is all algabra

OpenStudy (anonymous):

oh ok. so it's like: 2(a+h)^2+14(a)-13 over h?

OpenStudy (anonymous):

which comes to...

OpenStudy (anonymous):

no hold on

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

you have to subtract first

OpenStudy (anonymous):

well, first you have to expand, then you have to subtract

OpenStudy (anonymous):

\(g(a+h)= 2(a+h)^2+14(a+h)-13\)

OpenStudy (anonymous):

expand like 2a+2h.... and so on?

OpenStudy (anonymous):

(2a+2h)^2?

OpenStudy (anonymous):

for the first one

OpenStudy (anonymous):

yeah it is \[2(a^2+2ah+h^2)+14a+14h-13\] \[=2a^2+4ah+2h^2+14a+14h-13\]

OpenStudy (anonymous):

square first multiply by 2 second

OpenStudy (anonymous):

oh ok, gotcha

OpenStudy (anonymous):

from this you need to subtract \(2a^2+14a-13\) so everything without an \(h\) will be gone

OpenStudy (anonymous):

then you can factor out an \(h\) from the remaining terms in the numerator, and cancel with the \(h\) in the denominator

OpenStudy (anonymous):

you should be left with \[4ah +14h+2h^2\] before you divide after you divide you get \[4a+14\] and next week you will be able to do this problem instantly in your head

OpenStudy (anonymous):

ok, so I have 2a^2 +4ah + 2h^2 +14a +14h -13 and then I subtract 2a2+14a−13? how do I know to do that?

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

\(2a^2-2a^2=0\) \(14a-14a=0\) and \(-13-(-13)=0\)

OpenStudy (anonymous):

in other words, every term without an \(h\) is gone

OpenStudy (anonymous):

oh, now I get it

OpenStudy (anonymous):

and that leaves me with 4ah + 2h^2 +14h

OpenStudy (anonymous):

yes, and now divide by \(h\)

OpenStudy (anonymous):

as you said lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4a+2h+14

OpenStudy (anonymous):

which is one of my multiple choice answers...

OpenStudy (anonymous):

well, I dunno if I'll be able to do that in my head instantly, but if I can work it out, then I'll be ok at least haha. Thanks @satellite73